OpenStudy (anonymous):

In a computer catalog, a computer monitor is said to be 19 inches. This distance is the diagonal distance across the screen. If the screen is 10 inches high, what is the width of the screen? Round to the nearest tenth.

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

pythagorean theorem a^2 + b^2 = c^2 x^2 + 10^2 = 19^2 x^2 + 100 = 361 I'll let you finish

OpenStudy (anonymous):

multiply?

jimthompson5910 (jim_thompson5910):

in a way, but that's not the full story